OpenStudy (anonymous):

Henry and Layla are ordering flowers for a birthday party. Large arrangements are $75 and small arrangements are $50. They can afford to spend no more than $700, but would like to purchase more than 10 arrangements. Which is not a combination of large and small arrangements that Henry and Layla could purchase? 3 large and 8 small arrangements 4 large and 7 small arrangements 5 large and 6 small arrangements 7 large and 4 small arrangements

OpenStudy (anonymous):

5 large and 6 small arrangements.

OpenStudy (anonymous):

375 + 300 = 675.

OpenStudy (anonymous):

damm bro your amazing ! @Mikeyy1992

OpenStudy (anonymous):

Thanks, become my fan. I'll become yours and I can try to help you with any other questions.

jimthompson5910 (jim_thompson5910):

The answer of "5 large and 6 small arrangements" is incorrect

OpenStudy (anonymous):

@jim_thompson5910 how ? :c

jimthompson5910 (jim_thompson5910):

Mikeyy1992 just showed why it's not the answer

jimthompson5910 (jim_thompson5910):

the result (675) is less than 700, so that works they want something that doesn't work

jimthompson5910 (jim_thompson5910):

It says Which is NOT a combination of large and small arrangements that Henry and Layla could purchase?

OpenStudy (anonymous):

Can you show me how work this problem

jimthompson5910 (jim_thompson5910):

you just plug each point into 75x+50y and see which one gives a result over 700

OpenStudy (anonymous):

for example like 75(5)+50(6) ?

jimthompson5910 (jim_thompson5910):

yep, 75x+50y = 75*5+50*6 = 675 ... works so (5,6) is out

OpenStudy (anonymous):

ooooh lol :D